In a two-tailed test, a statistician got a z test statistic of 1.82. What is the p-value?

Multiple Choice

  • .0688

  • .0301

  • .0874

  • .0708
Answer #1

solution

this is two tailed test

P(z > 1.82) = 1 - P(z < 1.82) = 1-0.9656=0.0344*2=0.0688

P-value = 0.0688 ( using z table)